# VAMO: Efficient Large-Scale Nonconvex Optimization via Adaptive Zeroth Order Variance Reduction

### v1

@@ 3.4 Notations
In this paper, we denote $\nabla f(x),\nabla f_{i}(x)$ as first-order gradients of $f(x)$ and $f_{i}(x)$, respectively. Also denote ^$\hat{\nabla}f(x),\hat{\nabla}f_{i}(x)$ as their zeroth-order variants. $\mathbb{E}[\cdot]$ operates as the usual mathematical expectation, and $\mathcal{I}$ is a mini-batch of indices sampled from $[n]:={1,\dots,n}$, with size $b=|\mathcal{I}|$. $\|\cdot\|_{2}$ denotes the Euclidean/l2 norm as per usual.
@@ 4.1 From SVRG and ZO-SVRG to Hybrid SVRG
The principles of FO-SVRG and ZO-SVRG have been extensively explored in optimization literature reddi2016stochastic ; johnson2013accelerating ; liu2018zeroth ; ji2019improved . FO-SVRG, in particular, is known to achieve a linear convergence rate $O(1/T)$ for non-convex problems under certain conditions, significantly outperforming the convergence rate of FO-SGD reddi2016stochastic . The key step of FO-SVRG involves leveraging a full gradient $\nabla f(\hat{\mathbf{x}})$, computed periodically at a checkpoint $\hat{\mathbf{x}}$, to construct a variance-reduced stochastic gradient estimate johnson2013accelerating :
$\hat{\mathbf{g}}_{\text{FO-SVRG}}=\nabla f_{\mathcal{I}}(\mathbf{x})-\nabla f_{\mathcal{I}}(\hat{\mathbf{x}})+\nabla f(\hat{\mathbf{x}}),$ (4)
where $\nabla f_{\mathcal{I}}(\mathbf{x})=\frac{1}{b}\sum_{i\in\mathcal{I}}\nabla f_{i}(\mathbf{x})$ is the mini-batch stochastic gradient from a subset $\mathcal{I}\subseteq[n]$ of size $b$. A crucial property of $\hat{\mathbf{g}}_{\text{FO-SVRG}}$ is that $\hat{\mathbf{g}}_{\text{FO-SVRG}}$ is an unbiased gradient estimate of $\nabla f(\mathbf{x})$, $\mathbb{E}[\hat{\mathbf{g}}_{\text{FO-SVRG}}]=\nabla f(\mathbf{x})$.
@@
$\hat{\mathbf{g}}_{\text{ZO-SVRG}}=\hat{\nabla}f_{\mathcal{I}}(\mathbf{x})-\hat{\nabla}f_{\mathcal{I}_{k}}(\hat{\mathbf{x}})+\hat{\nabla}f(\hat{\mathbf{x}}),$ (5)
where ^$\hat{\nabla}f_{\mathcal{I}}(\mathbf{x})=(1/b)\sum_{i\in\mathcal{I}}\hat{\nabla}f_{i}(\mathbf{x})$, ^$\hat{\nabla}f(\mathbf{x})=\hat{\nabla}f_{[n]}(\mathbf{x})$, and ^$\hat{\nabla}f_{i}(\mathbf{x})$ is a ZO gradient estimate (e.g., two-point or multi-point as defined in Section 3.3). While structurally similar, a key distinction is that $\hat{\mathbf{g}}_{\text{ZO-SVRG}}$ is generally a biased estimate of $\nabla f(\mathbf{x})$ due to the inherent bias of ^$\hat{\nabla}f_{i}(\mathbf{x})$ relative to $\nabla f_{i}(\mathbf{x})$. This bias significantly complicates its convergence analysis compared to FO-SVRG.
VAMO (Algorithm 1) is motivated by the high cost of computing the full gradient $\nabla f(\hat{\mathbf{x}})$ in SVRG for large-scale models, and introduces a hybrid gradient estimator that combines FO and ZO components to reduce this overhead:
$\hat{\mathbf{g}}=\nabla f_{\mathcal{I}}(\mathbf{x})-\alpha\left(\hat{\nabla}f_{\mathcal{I}}(\hat{\mathbf{x}})-\hat{\nabla}f(\hat{\mathbf{x}})\right).$ (6)
Here, $\nabla f_{\mathcal{I}}(\mathbf{x})$ is the standard FO mini-batch stochastic gradient at the current iterate $\mathbf{x}$, while ^$\hat{\nabla}f(\hat{\mathbf{x}})$ is the ZO estimate of the full gradient at the checkpoint $\hat{\mathbf{x}}$.
A critical design choice in Eq. (6) is the construction of the variance-reduction term $\alpha\left(\hat{\nabla}f_{\mathcal{I}}(\hat{\mathbf{x}})-\hat{\nabla}f(\hat{\mathbf{x}})\right)$. To preserve the desirable unbiased property of SVRG, we ensure that the expectation of the ZO-based correction term, $\mathbb{E}\left[\hat{\nabla}f_{\mathcal{I}}(\hat{\mathbf{x}})-\hat{\nabla}f(\hat{\mathbf{x}})\right]$, is zero. Using ZO estimates for both terms within the parentheses, ^$\hat{\nabla}f_{\mathcal{I}}(\hat{\mathbf{x}})$ and ^$\hat{\nabla}f(\hat{\mathbf{x}})$, rather than mixing FO and ZO estimates within that difference, is key to this property and also contributes to computational savings at the checkpoint. Maintaining this unbiasedness is pivotal, as it allows for a more tractable convergence analysis akin to FO-SVRG, distinguishing our approach from many ZO algorithms that contend with biased estimators.
Furthermore, VAMO introduces a novel mixing coefficient $\alpha>0$. This parameter, absent in traditional SVRG or ZO-SVRG, allows for explicit control over the influence of the ZO-derived variance term. The choice of $\alpha$, which we will discuss in Section 5.1, enables a flexible balance between computational overhead and convergence performance.
The introduction of this hybrid structure, particularly the ZO estimation at checkpoints and the mixing coefficient $\alpha$, means that the convergence analysis of VAMO cannot be trivially inherited from existing FO-SVRG or ZO-SVRG analyses. It requires a dedicated theoretical investigation to characterize its behavior and prove its convergence guarantees, which constitutes a core part of our contribution in Section 4.2. This distinct analytical challenge underscores the theoretical novelty of our work.
Algorithm 1 VAMO$(T,m,\{\eta_{k}\},b,\bar{x}_{0},\mu,\alpha)$
@@
3:     compute ZO estimate $\hat{g}_{s}=\hat{\nabla}f(\bar{x}_{s-1})$
4:     set $x_{0}^{s}=\bar{x}_{s-1}$
@@
7:        compute hybrid FO and ZO gradient blending (6): $\mathbf{v}_{k}^{s}=\nabla f_{I_{k}}(x_{k}^{s})-\alpha(\hat{\nabla}f_{I_{k}}(x_{0}^{s})-\hat{g}_{s})$
8:        update $x_{k+1}^{s}=x_{k}^{s}-\eta_{k}\mathbf{v}_{k}^{s}$
@@
12:  return $\bar{x}$ chosen uniformly at random from $\{\{x_{k}^{s}\}^{m-1}_{k=0}\}^{S}_{s=1}$.
@@ 4.2 Convergence analysis
In this section, we present the convergence analysis for VAMO using the two-point ZO gradient estimate (Eq. (2)). Our analysis is based on an upper bound on the expected squared gradient norm $\mathbb{E}\bigl{[}\|\nabla f(\bar{\mathbf{x}})\|_{2}^{2}\bigr{]}$, as shown in Theorem 1. As discussed in Section 3.2, for non-convex objectives, a small value of $\mathbb{E}\bigl{[}\|\nabla f(\bar{\mathbf{x}})\|_{2}^{2}\bigr{]}$ implies convergence to a stationary point.
@@
where $T=Sm$, $f^{*}=\min_{\mathbf{x}}f(\mathbf{x})$, $\bar{\gamma}=\min_{k\in[m]}\gamma_{k}$, and $\chi_{m}=\sum_{k=0}^{m-1}\chi_{k}$ with
$\gamma_{k}$ $=\biggl{(}1-\frac{c_{k+1}}{\beta_{k}}\biggr{)}\eta_{k}-4\biggl{(}\frac{L}{2}+c_{k+1}\biggr{)}\times\biggl{(}2\alpha^{2}-2\alpha+1+\frac{24d\delta_{n}}{b}\alpha^{2}\biggr{)}\eta_{k}^{2},$ (8)
$\chi_{k}$ $=\biggl{(}\frac{L}{2}+c_{k+1}\biggr{)}\times\biggl{(}\frac{9\delta_{n}}{b}d^{2}L^{2}\mu^{2}\alpha^{2}+\frac{4\sigma^{2}}{b}\bigl{(}24d\delta_{n}\alpha^{2}+(1-\alpha)^{2}\bigr{)}\biggr{)}\eta_{k}^{2}.$ (9)
The coefficients $\{c_{k}\}$ are given by:
$c_{k}=\biggl{[}1+\beta_{k}\eta_{k}+\frac{6(4d+1)L^{2}\delta_{n}\eta_{k}^{2}}{b}\biggr{]}c_{k+1}+\frac{3(4d+1)L^{3}\delta_{n}\eta_{k}^{2}}{b},\quad c_{m}=0.$ (10)
Proof. See Appendix A.3.
Compared to the convergence rate of SVRG reddi2016stochastic , Theorem 1 has an additional error $\left(\frac{S\chi_{m}}{T\bar{\gamma}}\right)$ because of the use of ZO gradient estimator. $\chi_{m}$ depends on the epoch $m$, the step size $\eta_{k}$, the smoothing parameter $\mu$, the mini-batch size $b$, the number of optimization variables $d$ and the mixing constant $\alpha$. To obtain a clear dependence on these parameters and explore deeper convergence insights, we simplify (7) to suit specific parameter settings, as shown below.
@@
Proof. See Appendix A.4.
From Corollary 1, we can observe that one advantage of this algorithm is that, compared to previous ZO algorithms, the value of smoothing parameters $\mu$ is less restrictive. For example, ZO-SVRG liu2018zeroth required $\mu\leq O(d^{-1/2}T^{-1/2})$, and ZO-SGD ghadimi2013stochastic required $\mu\leq O(d^{-1}T^{-1/2})$. Compared to FO-SGD, the algorithm achieves an improved rate of $O(1/T)$ rather than the rate of $O(1/\sqrt{T})$. Compared to ZO algorithms, the convergence rate is independent of the number of optimization variables $d$. Compared to first-order SVRG. though both methods achieve a linear convergence rate, VAMO suffers an additional error of $O(1/b)$ inherited from $\frac{S\chi_{m}}{T\bar{\gamma}}$ in (1).
@@ 5 VAMO with Multi-Point ZO Gradient Estimation
Suppose assumptions A1 and A2 hold, and the multi-point ZO gradient estimate is used in Algorithm 1. The gradient norm bound in (7) yields the simplified convergence rate:
@@
With parameter choices $\mu=\frac{1}{q\sqrt{T}}$, $\eta=\frac{\rho}{L}$, $\alpha=\frac{q}{d}$, and $\beta=L$, the coefficients satisfy:
$c_{k}$ $=\biggl{(}1+\beta_{k}\eta_{k}\biggr{)}c_{k+1}+\biggl{(}\frac{L}{2}+c_{k+1}\biggr{)}\biggl{(}1+\frac{4d}{q}\biggr{)}\frac{12L^{2}\delta_{n}\eta_{k}^{2}\alpha^{2}}{b},$ (15)
$\gamma_{k}$ $=\eta_{k}-\frac{c_{k+1}\eta_{k}}{\beta_{k}}-4\biggl{(}\frac{L}{2}\eta_{k}^{2}+c_{k+1}\eta_{k}^{2}\biggr{)}\biggl{(}\Bigl{(}2+\frac{24d\delta_{n}}{qb}\Bigr{)}\alpha^{2}-2\alpha+1\biggr{)},$ (16)
$\chi_{k}$ $=\biggl{(}\frac{L}{2}\eta_{k}^{2}+c_{k+1}\eta_{k}^{2}\biggr{)}\times\biggl{(}\frac{3\delta_{n}}{b}\Bigl{(}1+\frac{2}{q}\Bigr{)}L^{2}d^{2}\mu^{2}\alpha^{2}+\frac{4\sigma^{2}}{b}\Bigl{(}\frac{24d\delta_{n}\alpha^{2}}{q}+(1-\alpha)^{2}\Bigr{)}\biggr{)}.$ (17)
Proof. See Appendix A.5.
By contrast with Corollary 1, it can be seen from Eq. (14) that the use of multi-point version of VAMO reduces the error $O(1/b)$ in Eq. (1) by leveraging multiple $q$ direction sampling, while increasing the computational cost accordingly. If $q=d$, the algorithm’s computational cost and convergence rate become comparable to FO-SVRG. Please note that the smoothing parameter $\mu$ is more restrictive than that in two-point version of VAMO for reducing the error. A comprehensive summary and comparison of the computational complexities and convergence rates of our proposed VAMO methods against various FO and ZO algorithms can be found in Table 1, which is presented and discussed in Section 2.
5.1 Balancing FO and ZO Information via $\alpha$
The mixing coefficient $\alpha$ in the VAMO update (Eq. (6)) critically balances the FO stochastic gradient $\nabla f_{\mathcal{I}}(\mathbf{x})$ against the ZO variance correction term ^$\hat{\nabla}f_{\mathcal{I}}(\hat{\mathbf{x}})-\hat{\nabla}f(\hat{\mathbf{x}})$. The optimal choice for $\alpha$ directly depends on the estimation error $\omega_{i}$ inherent in the ZO gradient components (^$\hat{\nabla}f_{i}(\mathbf{x})=\nabla f_{i}(\mathbf{x})+\omega_{i}$). As established in the literature liu2018zeroth ; liu2020primer and detailed in Appendix A.7, the expected squared ZO error $\mathbb{E}[|\omega_{i}|_{2}^{2}]$ typically scales as $\mathcal{O}(d)$ for two-point estimates and $\mathcal{O}(d/q)$ for multi-point estimates using $q$ random directions. fThis relationship dictates that $\alpha$ should reflect the trustworthiness of the ZO estimates: when the ZO error is substantial (e.g., large $d$, small $q$), a smaller $\alpha$ is warranted to prevent amplifying this error. Conversely, when ZO estimates are more reliable (e.g., larger $q$ reducing error), a larger $\alpha$ can more aggressively leverage the variance reduction. This principled inverse relationship between ZO error magnitude (influenced by $d$ and $q$) and the appropriate scale of $\alpha$ is key. While specific forms like $\alpha\propto 1/d$ or $\alpha\propto q/d$ analyzed in our theoretical sections (e.g., Corollary 1 and Theorem 2) illustrate this adaptive trend, the core insight is that $\alpha$ must be adjusted to counterbalance the ZO estimator’s error profile. Such adaptability enables VAMO to effectively navigate the trade-off between computational cost and convergence performance, a central aspect of its practical utility.
@@ 6 Applications and Experiments
In this section, we present empirical results to validate the effectiveness and adaptability of the proposed VAMO. We first demonstrate VAMO’s adaptive capabilities by evaluating the impact of its key tunable component, the number of ZO random directions ($q$), on a synthetic task, showcasing how performance can be configured based on computational budget. Subsequently, we benchmark its performance against standard FO and ZO methods on a classification task using DNNs, and finally, we demonstrate its utility in a large model fine-tuning scenario gautam2024variance . Detailed experimental setups and hyperparameter settings for all methods and tasks are provided in Appendix A.8,  A.9 and  A.10.
Adaptability Experiment: Nonconvex Least Squares with Varying $q$ To showcase VAMO’s adaptive nature and empirically validate the theoretical benefits of its multi-point ZO estimation strategy (as discussed in Section 5), we conducted experiments on a synthetic non-convex least-squares task. The objective was $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}(h(\mathbf{x};\mathbf{z}_{i})-y_{i})^{2}$, with $n=1000$ component functions and a parameter dimension $d=100$, where $h(\mathbf{x};\cdot)$ was parameterized by a simple non-convex neural network. We compared VAMO variants using $q\in\{1,3,5\}$ (number of ZO query directions for the multi-point ZO gradient estimator) against the classical FO-SGD algorithm.
Fig. 1(a) presents the training loss convergence. Consistent with our theoretical analysis (see Table 1), all VAMO variants achieve an $O(1/T)$ convergence rate, outperforming FO-SGD’s $O(1/\sqrt{T})$ rate. The figure clearly illustrates VAMO’s adaptability: increasing $q$ improves convergence performance, effectively mitigating the additional $O(1/b)$ error term associated with the two-point ($q=1$) variant. This aligns with the theoretical prediction that this error term diminishes for larger $q$ (scaling towards $O(\frac{1}{b}(1-q/d)^{2})$). These results empirically validate our theory and highlight VAMO’s practical ability to adaptively trade computational cost (via varying $q$) for enhanced convergence by managing ZO estimation error, a key aspect of its flexible and adaptive design.
Multiclass Classification For this benchmark on the MNIST dataset lecun1998gradient , we trained a Multi-Layer Perceptron (MLP) and compared our proposed VAMO (two-point version, $q=1$) against FO-SGD robbins1951stochastic , ZO-SGD ghadimi2013stochastic , and ZO-SVRG liu2018zeroth . As illustrated by the training loss convergence in Fig. 1(b), our VAMO algorithm demonstrates a significant performance advantage over the purely ZO methods (ZO-SGD and ZO-SVRG), achieving both substantially faster convergence and a better final loss value. Moreover, VAMO’s convergence behavior is highly competitive with that of the standard FO-SGD algorithm. These findings underscore the practical effectiveness of our hybrid strategy.
GPT2 Fine-Tuning To further assess VAMO’s practical utility and its advantages in complex, large-scale settings, we applied it to the task of fine-tuning a pre-trained GPT-2 model radford2019language . Specifically, this experiment involved fine-tuning the base GPT-2 model on the MultiNLI (MNLI) dataset williams2017broad for a natural language inference task. Our proposed VAMO algorithm was benchmarked against key FO optimizers, notably FO-SGD robbins1951stochastic , and representative ZO methods such as ZO-SGD and ZO-SVRG.
In Fig. 2, we present training loss against iteration steps and queries, highlighting VAMO’s practical advantages, with the query-based evaluation in Fig. 2(b) focusing on its comparison with FO-SGD. To clearly distinguish from potential ZO query definitions and to establish a consistent basis for comparison with FO methods, a query here is specifically defined in terms of FO computational units: it denotes a single pass (either forward or backward) of one data sample through the model; consequently, an FO mini-batch gradient calculation on $b$ samples costs $2b$ queries. According to Table 1, VAMO incurs an additional query cost over FO-SGD due to the use of the full ZO gradient. FO-SGD has a total computational cost of $O(bdT)$, while VAMO incurs an additional $O(nS)$ cost. For large models such as GPT-2 (e.g., $d\approx 1.2\times 10^{8},n\approx 256,m=10,b=32$, with $T=Sm$), this leads to only a minor fractional overhead of approximately $n/(2bdm)$ compared to FO-SGD, assuming the per-step ZO correction is also query-light. Therefore, with a total query cost only slightly higher than that of FO-SGD, VAMO achieves significantly faster and more stable convergence per query, as demonstrated in Fig. 2(b). This improvement stems from its variance-reduced SVRG backbone, now made efficient and practical through ZO techniques.
@@
(a) Training loss versus iteration steps
(b) Training loss versus queries
Figure 2: Convergence comparison on the fine-tuning GPT2 task against pure FO and ZO methods.

### v2

@@ 3.4 Notations
In this paper, we denote $\nabla f(x),\nabla f_{i}(x)$ as FO gradients of $f(x)$ and $f_{i}(x)$, respectively and $\hat{\nabla}f(x),\hat{\nabla}f_{i}(x)$ as their ZO variants. $\mathbb{E}[\cdot]$ operates as the usual mathematical expectation, and $\mathcal{I}$ is a mini-batch of indices sampled from $n]:={1,\dots,n}$, with size $b=|\mathcal{I}|$. $\|\cdot\|_{2}$ denotes the Euclidean (i.e., $\ell_{2}$) norm.
@@ 4.1 From SVRG and ZO-SVRG to Hybrid SVRG
[…]
where $\nabla f_{\mathcal{I}}(\mathbf{x})=\frac{1}{b}\sum_{i\in\mathcal{I}}\nabla f_{i}(\mathbf{x})$ is the mini-batch stochastic gradient from a subset $\mathcal{I}\subseteq[n]$ of size $b$. A crucial property of $\hat{\mathbf{g}}_{\text{FO-SVRG}}$ is that $\hat{\mathbf{g}}_{\text{FO-SVRG}}$ is an unbiased gradient estimate of $\nabla f(\mathbf{x})$, $\mathbb{E}[\hat{\mathbf{g}}_{\text{FO-SVRG}}]=\nabla f(\mathbf{x})$ (Johnson & Zhang, 2013).
@@
[…]
where $\hat{\nabla}f_{\mathcal{I}}(\mathbf{x})=(1/b)\sum_{i\in\mathcal{I}}\hat{\nabla}f_{i}(\mathbf{x})$, $\hat{\nabla}f(\mathbf{x})=\hat{\nabla}f_{[n]}(\mathbf{x})$, and $\hat{\nabla}f_{i}(\mathbf{x})$ is a ZO gradient estimate, as defined in Section 3.3. While structurally similar, a key distinction is that $\hat{\mathbf{g}}_{\text{ZO-SVRG}}$ is generally a biased estimate of $\nabla f(\mathbf{x})$ due to the inherent bias of $\hat{\nabla}f_{i}(\mathbf{x})$ relative to $\nabla f_{i}(\mathbf{x})$ (Liu et al., 2020). This bias significantly complicates its convergence analysis compared to FO-SVRG.
[…]
The introduction of this hybrid structure, particularly the ZO estimation at snapshot and the mixing coefficient $\alpha$, means that the convergence analysis of VAMO cannot be trivially inherited from existing FO-SVRG or ZO-SVRG analyses. It requires a dedicated theoretical investigation to characterize its behavior and prove its convergence guarantees, which constitutes a core part of our contribution in Section 4.2. This distinct analytical challenge underscores the theoretical novelty of our work.
Algorithm 1 VAMO $(T,m,\{\eta_{k}\},b,\bar{\mathbf{x}}_{0},\mu,\alpha)$
@@
3:  compute ZO estimate $\hat{\mathbf{g}}_{s}=\hat{\nabla}f(\bar{\mathbf{x}}_{s-1})$
4:  set $\mathbf{x}_{0}^{s}=\bar{\mathbf{x}}_{s-1}$
@@
7:   compute hybrid FO and ZO gradient blending: $\mathbf{v}_{k}^{s}=\nabla f_{I_{k}}(\mathbf{x}_{k}^{s})-\alpha(\hat{\nabla}f_{I_{k}}(\mathbf{x}_{0}^{s})-\hat{\mathbf{g}}_{s})$
8:   update $\mathbf{x}_{k+1}^{s}=\mathbf{x}_{k}^{s}-\eta_{k}\mathbf{v}_{k}^{s}$
@@
12: return $\bar{\mathbf{x}}$ chosen uniformly at random from $\{\{\mathbf{x}_{k}^{s}\}^{m-1}_{k=0}\}^{S}_{s=1}$.
@@ 4.2 Convergence analysis
In this section, we present the convergence analysis for VAMO using the two-point ZO gradient estimate in  equation 2. Our analysis is based on an upper bound on the expected squared gradient norm $\mathbb{E}\bigl[\|\nabla f(\bar{\mathbf{x}})\|_{2}^{2}\bigr]$, as shown in Theorem 1. As discussed in Section 3.2, for non-convex objectives, a small value of $\mathbb{E}\bigl[\|\nabla f(\bar{\mathbf{x}})\|_{2}^{2}\bigr]$ implies convergence to a stationary point.
@@
where $T=Sm$, $f^{*}=\min_{\mathbf{x}}f(\mathbf{x})$, $\bar{\gamma}=\min_{k\in[m]}\gamma_{k}$, and $\chi_{m}=\sum_{k=0}^{m-1}\chi_{k}$. $\gamma_{k}$ and $\chi_{k}$ are coefficients which depend on $\{\eta_{k},\mu,b,d,\alpha\}$. The proof is provided in Appendix E.
Compared to FO-SVRG, Theorem 1 has an additional error $(S\chi_{m}/(T\bar{\gamma}))$ due to the use of the ZO gradient estimator. To obtain a clear dependence on these parameters and explore deeper insights into convergence, we simplify equation 7 to suit the specific parameter settings, as shown below.
@@
The proof is provided in Appendix F.
From Corollary 1, we can observe that one advantage of VAMO is that, compared to previous ZO algorithms, the value of smoothing parameters $\mu$ is less restrictive. For example, ZO-SVRG required $\mu\leq O(1/\sqrt{dT})$, and ZO-SGD required $\mu\leq O(1/d\sqrt{T})$ (Liu et al., 2018b). Compared to FO-SGD, the algorithm achieves an improved rate of $O(1/T)$ rather than the rate of $O(1/\sqrt{T})$. Compared to ZO algorithms, the convergence rate is independent of the parameter dimension $d$. Compared to FO-SVRG, VAMO suffers an additional error of $O(1/b)$ inherited from $(S\chi_{m}/(T\bar{\gamma}))$ in equation 1.
4.3 Memory Efficiency of VAMO
Building on the convergence analysis of VAMO in the Section 4.2, where we showed that VAMO achieves an $O(1/T)$ convergence rate, we now analyze its memory efficiency in large-scale training.
In large-scale training, the total memory footprint of an algorithm is dominated by three components: model parameters $|\mathbf{x}|$, optimizer states, and dynamic memory for gradient computation (Zhang et al., 2024). Here $|\mathbf{x}|$ denotes the memory required to store the model parameters in full precision, and $|\mathbf{x}_{l}|$ is that of the parameters in layer $l$. The symbol $|\mathbf{a}_{l}|$ denotes the memory of activations or intermediate results of one single sample at layer $l$.
For FO methods, dynamic memory is dominated by intermediate results, scaling roughly as $\sum_{l}b\cdot|\mathbf{a}_{l}|$ in Table 1 for a mini-batch of size $b$. FO-SVRG (Johnson & Zhang, 2013) further requires computing the full gradient at the snapshot point; while this can be accumulated over smaller mini-batches to reduce peak memory of intermediate results to $\sum_{l}b\cdot|\mathbf{a}_{l}|$, it increases computational cost per epoch due to multiple passes over all $n$ samples. If the full FO gradient is computed at once, the dynamic memory of intermediate results will increase to $\sum_{l}n\cdot|\mathbf{a}_{l}|$, which poses a substantial memory burden for large-scale settings. Therefore, FO-SVRG must trade off compute and memory when computing full FO gradients. In contrast, VAMO leverages full ZO gradient estimates, which do not require backpropagation, and thus do not need to store intermediate results. Even for estimation of full batch gradients at once, dynamic memory only reaches $\max_{l}|\mathbf{x}_{l}|$ , independent of $b$ , the same as ZO-SGD, which is much smaller than $\sum_{l}b\cdot|\mathbf{a}_{l}|$. Therefore, VAMO do not need to introduce a compute–memory trade-off like FO-SVRG in large-scale settings. Regarding optimizer states, VAMO only needs $|\mathbf{x}|$ for storing full ZO gradient at the snapshot, substantially smaller than Adagrad and Adam which need around $2|\mathbf{x}|$ or $3|\mathbf{x}|$ to store first- or second-momentum. In training large models, the optimizer states of Adam (Kingma & Ba, 2014) and Adagrad (Duchi et al., 2011) can occupy a substantial portion of memory, posing a significant challenge in resource-constrained environments. Table 3 summarizes the three memory components for different optimizers and a more detailed theoretical and empirical analysis of memory efficiency is provided in Appendix A.
@@ 5 VAMO with Multi-Point ZO Gradient Estimation
Suppose assumptions A1 and A2 hold, and the multi-point ZO gradient estimate is used in Algorithm 1. The gradient norm bound in equation 7 yields the simplified convergence rate:
@@
With parameter choices $\mu=\frac{1}{q\sqrt{T}}$, $\eta=\frac{\rho}{L}$, $\alpha=\frac{q}{d}$. The proof is provided in Appendix G.
By contrast with Corollary 1, it can be seen from  equation 11 that the use of multi-point version of VAMO reduces the error $O(1/b)$ in  equation 1 by leveraging multiple $q$ direction sampling, while increasing the computational cost accordingly. If $q=d$, the algorithm’s convergence rate become comparable to FO-SVRG. A comprehensive summary and comparison of the computational complexities and convergence rates of our proposed VAMO methods against various FO and ZO algorithms can be found in Table 1 presented in Section 2. Briefly, the mixing coefficient $\alpha$ governs the trade-off between the FO stochastic gradient and the ZO variance-correction: larger $\alpha$ increases reliance on ZO information while smaller $\alpha$ favors FO updates to mitigate the error of ZO estimation. A detailed theoretical and empirical analysis of $\alpha$ is provided in Appendix J.2.
@@ 6 Applications and Experiments
In this section, we present empirical results to validate the effectiveness and adaptability of the proposed VAMO through three experiments: (i) an adaptability study on a synthetic task varying the number of ZO query directions $q$, (ii) a benchmark on MNIST classification (LeCun et al., 1998), and (iii) large-scale fine-tuning of GPT-2, GPT-2 (Radford et al., 2019) and RoBERTa (Liu et al., 2019) on SST-2 (Socher et al., 2013) and MNLI (Williams et al., 2017). We benchmark against popular FO methods (FO-SGD (Robbins & Monro, 1951), FO-Adagrad (Duchi et al., 2011), FO-Adam (Kingma & Ba, 2014)) and ZO methods (ZO-SGD (Ghadimi & Lan, 2013), ZO-SVRG (Liu et al., 2018b)). Full setups and hyperparameters are deferred to Appendix I.
@@
Adaptability Experiment.
To highlight VAMO’s adaptive nature, we tested its multi-point ZO estimation strategy on a synthetic non-convex least-squares task. We compared variants using $q\in\{1,3,5\}$ query directions against FO-SGD. Fig. 1(a) presents the training loss convergence. Consistent with our theoretical analysis in Table 1, all VAMO variants achieve an $O(1/T)$ convergence rate, outperforming FO-SGD’s $O(1/\sqrt{T})$ rate. The figure clearly illustrates VAMO’s adaptability: increasing $q$ improves convergence performance, effectively mitigating the additional $O(1/b)$ error term associated with the two-point ($q=1$) variant. This aligns with the theoretical prediction that this error term diminishes for larger $q$ (scaling towards $O((1-q/d)^{2}/b$). These results empirically validate our theory and highlight VAMO’s practical ability to adaptively trade computational cost for enhanced convergence by managing ZO estimation error, a key aspect of its adaptive design.
Multiclass Classification.
On the MNIST benchmark, we trained a Multi-Layer Perceptron (MLP) and compared VAMO (two-point version, $q=1$) against FO optimizer (FO-SGD) and ZO optimizers (ZO-SGD and ZO-SVRG). As shown in Fig. 1(b), VAMO significantly outperforms purely ZO methods, converging faster and reaching a lower final loss. Its performance is also highly competitive with FO-SGD, underscoring the practical effectiveness of our hybrid strategy.
Table 2: We measure peak GPU memory consumption (in GB) when fine-tuning RoBERTa-Large and GPT-2 Medium models under varying batch sizes (bs) with a fixed context length (cl=128).
RoBERTa-Large GPT-2-Medium
Method bs = 16 bs = 32 bs = 64 bs = 16 bs = 32 bs = 64
FO-SGD 4.63 6.72 10.83 7.08 11.50 20.33
FO-Adagrad 7.21 10.46 16.54 10.25 15.74 27.57
FO-Adam 8.62 11.59 17.66 11.59 17.05 29.07
VAMO 5.95 7.83 11.96 8.40 12.62 21.46
Fine-Tuning Experiments.
To further assess VAMO’s practical utility and its advantages in complex, large-scale settings, we further evaluate VAMO on fine-tuning pre-trained large models, including GPT-2, GPT-2 Medium, and RoBERTa-large, on the SST-2 and MNLI datasets, comparing against standard FO optimizers (FO-SGD, FO-Adagrad, FO-Adam) and ZO optimizers (ZO-SGD, ZO-SVRG). Table 2 presents the peak GPU memory usage of different algorithms for batch sizes of 16, 32, and 64, with the context length fixed at 128. Fig. 2 presents the training loss of different algorithms over both iteration steps and wall-clock time. VAMO achieves markedly faster and more stable convergence than FO-SGD, while maintaining a similar memory footprint. This improvement is theoretically supported by Table 1: VAMO enjoys a convergence rate of $\mathcal{O}(1/T)$, superior to the $\mathcal{O}(1/\sqrt{T})$ rate of FO-SGD. Compared with ZO methods, VAMO not only yields better empirical performance but also avoids dependence on the parameter dimension $d$, demonstrating stronger scalability to large models. When compared to Adagrad, VAMO achieves a similar level of performance but with far lower memory overhead, further validating its practicality. Although VAMO underperforms Adam in terms of empirical convergence speed, it remains much more memory-efficient, as shown in Table 2. Overall, these results confirm that VAMO provides a favorable balance between convergence efficiency and memory cost, making it a competitive optimizer for fine-tuning large models.
Figure 2: Convergence comparison on fine-tuning GPT-2, GPT-2 medium and RoBERTa-large on SST-2 and MNLI datasets against pure FO and ZO methods.